A sample of 36 randomly selected students has a mean test score of 83.4 with a standard deviation of 8.92. Assume the population has a normal distribution.

Find the margin of error, and then find the 95% confidence interval for the population mean.
Answer #1

df =n -1 = 36 - 1 = 35

From T table,

t critical value at 0.05 significance level with 35 df = 2.030

Margin of error = t * S / sqrt(n)

= 2.030 * 8.92 / sqrt( 36)

= 3.018

95% confidence interval for is

- E < < + E

83.4 - 3.018 < < 83.4 + 3.018

80.382 < < 86.418

95% CI is ( 80.382 , 86.418 )